An arithmetic student needs an average of 70 or more to receive credit for the course. She scored 82, 64, and 98 on the first three exams. Write an inequality representing the score she must get on the last test to receive credit for the course.
Answer by checkley71(8403)

ADDING THE FIRST THREE TESTS WE GET 244 THEREFORE SHE MUST GET AT LEAST X ON HER FOURTH TEST TO ACHIEVE AN AVERAGE OF 70
(244+X)/4=70 OR 244+X=280 OR X=280-244 OR X=36